Expected cash dividends are $2.00, the dividend yield is 8%, flotation costs are 6% of price, and the growth rate is 3%. Compute cost of new common stock. (Do not round intermediate calculations. Round your answer to 2 decimal places.) 13.61% 11.26% 12.51% 11.51%
Answer #1

D. 11.51%

Price = $2.00 / 0.08 = $25

Cost of new common stock = D1 / (P0 - F) + g

Cost of new common stock = [$2.00 / ($25 - (0.06 * $25))] + 0.03

Cost of new common stock = 0.0851 + 0.03

Cost of new common stock = 0.1151 or 11.51%